These word problems are killing me today. The value:v(in dollars per acre) of a wheat crop (d) days after planting is given by v=6d-.05d^2 for 20 Kyle
Answer by checkley77(12844)

v=6d-.05d^2 for d=20
v=6*20+.05*20^2
v=120+.05*400
v=120+20
v=140 dollars per acre after 20 days.